# Generative Enzyme Design Guided by Functionally Important Sites and Small-Molecule Substrates

### v3

@@ Abstract
Enzymes are genetically encoded biocatalysts capable of accelerating chemical reactions. How can we automatically design functional enzymes? In this paper, we propose EnzyGen, an approach to learn a unified model to design enzymes across all functional families. Our key idea is to generate an enzyme’s amino acid sequence and their three-dimensional (3D) coordinates based on functionally important sites and substrates corresponding to a desired catalytic function. These sites are automatically mined from enzyme databases. EnzyGen consists of a novel interleaving network of attention and neighborhood equivariant layers, which captures both long-range correlation in an entire protein sequence and local influence from nearest amino acids in 3D space. To learn the generative model, we devise a joint training objective, including a sequence generation loss, a position prediction loss and an enzyme-substrate interaction loss. We further construct EnzyBench, a dataset with 3157 enzyme families, covering all available enzymes within the protein data bank (PDB). Experimental results show that our EnzyGen consistently achieves the best performance across all 323 testing families, surpassing the best baseline by 10.79% in terms of substrate binding affinity. These findings demonstrate EnzyGen’s superior capability in designing well-folded and effective enzymes binding to specific substrates with high affinities. The code, model and dataset are released at https://github.com/LeiLiLab/EnzyGen.
@@ 1 Introduction
Enzymes are biological catalysts to accelerate challenging chemical reactions underlying a range of biological processes. They enjoy widespread applications in the production of pharmaceuticals (Wu et al., 2012), specialty chemicals (Carbonell et al., 2018) and biofuels (Liao et al., 2016). In enzymatic reactions, the substrate is a small molecule being converted by the enzyme catalyst. By binding with and acting on specific substrates, enzymes allow for dramatically accelerated rates in the transformation of their substrates (Bar-Even et al., 2011). Designing enzymes that can bind to specific substrates is a critical yet challenging problem.
@@
Recently, deep learning methods are promising in protein design (Huang et al., 2016; Pearce & Zhang, 2021). Existing deep learning approaches for functional protein design fall into three categories: (a) generative models for protein sequence design guided by fitness landscape (Brookes et al., 2019; Rives et al., 2021; Ren et al., 2022; Song & Li, 2023; Wang et al., 2023); (b) structure to sequence generation based on a targeted protein backbone (Ingraham et al., 2019; Dauparas et al., 2022; Zheng et al., 2023); (c) co-design of a backbone structure and a protein sequence that encodes it (Anishchenko et al., 2021; Wang et al., 2022; Shi et al., 2022; Yeh et al., 2023). However, these methods face many limitations on enzyme design. First, fitness-guided approaches are limited by the lack of fitness data for the vast majority of enzyme families. Second, the structures of many enzymes remain unknown (Binz et al., 2005; Fischman & Ofran, 2018). Third, prior approaches do not model substrates in enzyme design process. Finally, there is no unified model applicable to all enzyme families.
@@
In this paper, we aim to learn a unified generative model to design functional enzymes across thousands of enzyme families. The key design idea lies in the notion that the biological function of an enzyme is enabled by a subset of residues (amino acids), known as functionally important sites (Bickel et al., 2002; Chakrabarti & Lanczycki, 2007; Wang et al., 2022), and that an ideal enzyme should be able to bind its substrate(s) in enzymatic reactions. Therefore, we formulate the enzyme design problem as jointly generating the enzyme sequence and backbone structure given automatically mined functionally important sites and substrates related to specific enzyme functions.
@@
To this end, we propose EnzyGen, a unified model to co-design enzyme sequence and backbone structure. EnzyGen comprises an enzyme modeling module and a substrate representation module. We devise neighborhood attentive equivariant layers (NAELs) to jointly model enzyme amino acid sequence and residue coordinates. Each NAEL consists of a global attention sub-layer and a neighborhood equivariant sub-layer. The global attention sub-layer captures correlations among all residues within an entire enzyme sequence, while the neighborhood equivariant sub-layer updates residue representations and coordinates based on nearest neighbors in the 3D space. This architectural design facilitates information exchange with varying levels of granularity, promoting a comprehensive interaction among different residues. The substrate representation module also uses neighborhood equivariant layers to model a given substrate. Furthermore, We add category tags that correspond to BRENDA enzyme families to guide the design of enzymes with specific functions. To train EnzyGen, we design a joint training objective, which includes an amino acid type prediction loss, a coordinate reconstruction loss and an enzyme-substrate interaction loss.
@@
  • We propose EnzyGen to jointly design enzyme sequence and backbone structure. EnzyGen is the first unified enzyme design model across thousands of enzyme families.
  • We create EnzyBench, a comprehensive benchmark for enzyme design, including $101,974$ PDB entries across $3,157$ enzyme families, and a set of three metrics to evaluate enzyme quality before wet-lab experiments. EnzyBench includes a test set with $323$ families for various function validation.
  • We conduct experiments to evaluate EnzyGen and prior functional protein design methods. These results show that enzymes designed by EnzyGen lead to the best average enzyme-substrate ESP score (Kroll et al., 2023) of 0.65, substrate binding affinity of -9.44, and AlphaFold2 pLDDT (Jumper et al., 2021) of 87.45 across 323 families (BRENDA enzyme classification fourth-level categories). EnzyGen outperforms the best baseline by 10.79% in terms of substrate binding affinity.
@@ 2 Related Work
Methods for Functional Protein Design. Functional protein design has been studied with a wide variety of methods. Some works focus on designing functional proteins guided by fitness landscape, including searching (Brookes & Listgarten, 2018; Brookes et al., 2019; Kumar & Levine, 2020; Das et al., 2021; Hoffman et al., 2022; Melnyk et al., 2021; Anishchenko et al., 2021; Ren et al., 2022) or directly generating (Jain et al., 2022; Song & Li, 2023) protein sequences through deep generative models. Another category of methods concentrates on designing protein sequences capable of folding into backbone structures to fulfill specific functions (Wang et al., 2018; Strokach et al., 2020; Dauparas et al., 2022; Hsu et al., 2022; Sumida et al., 2023). Owing to the availability of large-scale data, some models are trained on extensive protein sequences from diverse families (Ferruz et al., 2022) or incorporate additional control tags for enhanced modeling (Madani et al., 2023). On the path from sequence to function, the role of structure emerges as a crucial intermediary. Consequently, some studies focus on the design of functional structures (Trippe et al., 2022; Watson et al., 2023). Recognizing the inherently complex interplay between protein sequence and structure, another set of works relies on the simultaneous generation of a new backbone structure and an amino acid sequence that folds into it (Anishchenko et al., 2021; Wang et al., 2022; Shi et al., 2022; Yeh et al., 2023; Song et al., 2023). Compared to the design of other functional proteins, functional enzyme design is largely underdeveloped. Some of the above methods have been applied to specific enzyme families, such as generating satisfactory enzyme sequences using deep generative models (Detlefsen et al., 2022; Lin et al., 2022a; Giessel et al., 2022). However, there is currently no unified generative model that is applicable to all enzyme families.
@@ 3 Proposed Method: EnzyGen
An enzyme consists of a chain of amino acids (also called residues) connected by peptide bonds, which folds into a proper 3D structure. Let $\mathcal{A}$ be the set of 20 common amino acids. We denote the sequence of a $N$-residue enzyme by $\boldsymbol{s}=\{s_{1},s_{2},...,s_{N}\}\in\mathcal{A}^{N}$ and their $C_{\alpha}$ coordinates by $\boldsymbol{x}=[\boldsymbol{x}_{1},\boldsymbol{x}_{2},...,\boldsymbol{x}_{N}]^{T}\in\mathbb{R}^{N\times 3}$. For residue type $s_{i}\in\mathcal{A}$, where $i\in\{1,2,...,N\}$, we denote its one-hot encoding as $\boldsymbol{s}_{i}=\text{onehot}(s_{i})$ and the functionally important site index set as $\mathcal{M}$. We denote a substrate as $\mathcal{V}=\{v_{1},...,v_{m}\}$, where $m$ is the number of its atoms, and $v_{j}=(\boldsymbol{h}_{j},\boldsymbol{x}_{j})$ denotes an atom where $j\in\{1,2,...,m\}$. Here, $\boldsymbol{h}_{j}$ represents pre-computed chemical features, and $\boldsymbol{x}_{j}\in\mathbb{R}^{3}$ is the corresponding coordinate. Each substrate is paired with a binary label $y$ indicating whether the enzyme can bind with it. Inspired by multilingual machine translation models in natural language processing, which employ language tags to guide model generation in specific target languages (Johnson et al., 2017; Liu et al., 2020), we similarly use enzyme family tags to facilitate enzyme design with desired functions. To provide a systematical and professional enzyme family classification, we leverage the enzyme classification (EC) tree in BRENDA (Schomburg et al., 2004) as illustrated in Appendix Figure 5.
The problem studied in this paper can be formulated as follows: given a desired enzyme family $c$ in EC tree, a substrate $\mathcal{V}$, a functionally important site index set $\mathcal{M}$, its residue set $\boldsymbol{s}_{\mathcal{M}}$ and the corresponding 3D coordinates $\boldsymbol{x}_{\mathcal{M}}$, generate an enzyme sequence $\boldsymbol{s}$ and $C_{\alpha}$ coordinates $\boldsymbol{x}$ of $N$ residues satisfying the substrate binding constraint $y$. Essentially, we aim to learn a generative model with probability $P(\boldsymbol{s},\boldsymbol{x},y|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c,\mathcal{V})$.
@@ 3.1 Overall Model Architecture
$\footnotesize\begin{split}P(\boldsymbol{s},\boldsymbol{x},y&|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c,\mathcal{V};\theta)=P(\boldsymbol{s},\boldsymbol{x}|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)\\ &\qquad\qquad\qquad\qquad\cdot P(y|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c,\mathcal{V};\theta)\\ P(\boldsymbol{s},\boldsymbol{x}&|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)=P(\boldsymbol{s}|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)\\ &\qquad\qquad\qquad\quad\cdot P(\boldsymbol{x}|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)\\ P(\boldsymbol{s}&|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)=\Pi_{i=1\&i\notin\mathcal{M}}^{N}P(\boldsymbol{s}_{i}|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)\\ P(\boldsymbol{s}_{i}&|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)=\mathrm{Softmax}(W_{\mathcal{A}}\cdot\boldsymbol{h}_{i}^{L_{e}})\\ P(\boldsymbol{x}&|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)=\Pi_{i=1\&i\notin\mathcal{M}}^{N}P(\boldsymbol{x}_{i}|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)\\ \boldsymbol{x}_{i}&|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}}\sim\mathcal{N}(\boldsymbol{x}_{i}^{L_{e}};\lambda I)\end{split}$ (1)
where $W_{\mathcal{A}}$ is the embedding matrix for 20 common amino acids, $\boldsymbol{h}_{i}^{L_{e}}$ is the output embedding for $i^{th}$ residue at the last layer of enzyme modeling module and $\boldsymbol{x}_{i}^{L_{e}}$ is the corresponding output coordinate. $\mathcal{N}(\boldsymbol{x}_{i}^{L_{e}};\lambda I)$ is the Gaussian distribution with mean $\boldsymbol{x}_{i}^{L_{e}}$ and covariance matrix $\lambda I$ ($I$ is the identity matrix). $\lambda$ is a hyperparameter. To find the optimal $\theta$, we maximize the conditional log likelihood (i.e., minimizing the negative log likelihood):
@@
$\small\log P(\boldsymbol{x}_{i}|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)=-\frac{\lambda}{2}||\boldsymbol{x}_{i}-\boldsymbol{x}_{i}^{L_{e}}||_{2}^{2}+\mathrm{const}$ (3)
@@
$\footnotesize\begin{split}&\mathcal{L}(\theta)=-\sum\nolimits_{i=1\&i\notin\mathcal{M}}^{N}\log P(\boldsymbol{s}_{i}|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)\\ &+\frac{\lambda}{2}\sum\nolimits_{i=1\&i\notin\mathcal{M}}^{N}||\boldsymbol{x}_{i}-\boldsymbol{x}_{i}^{L_{e}}||_{2}^{2}-\log P(y|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c,\mathcal{V};\theta)\end{split}$ (4)

### v4

@@ Abstract
Enzymes are genetically encoded biocatalysts capable of accelerating chemical reactions. How can we automatically design functional enzymes? In this paper, we propose EnzyGen, an approach to learn a unified model to design enzymes across massive functional families. Our key idea is to generate an enzyme’s amino acid sequence and their three-dimensional (3D) coordinates based on functionally important sites and substrates corresponding to a desired catalytic function. These sites are automatically mined from enzyme databases. EnzyGen consists of a novel interleaving network of attention and neighborhood equivariant layers, which captures both long-range correlation in an entire protein sequence and local influence from nearest amino acids in 3D space. To learn the generative model, we devise a joint training objective, including a sequence generation loss, a position prediction loss and an enzyme-substrate interaction loss. We further construct EnzyBench, a dataset with 3157 enzyme families, covering all available enzymes within the protein data bank (PDB). Experimental results show that our EnzyGen consistently achieves the best performance across all 323 testing families, surpassing the best baseline by 10.79% in terms of substrate binding affinity. These findings demonstrate EnzyGen’s superior capability in designing well-folded and effective enzymes binding to specific substrates with high affinities. The code, model and dataset are released at https://github.com/LeiLiLab/EnzyGen.
@@ 1 Introduction
Enzymes are biological catalysts to accelerate challenging chemical reactions underlying a range of biological processes. They enjoy widespread applications in the production of pharmaceuticals (Wu et al., 2012), specialty chemicals (Carbonell et al., 2018) and biofuels (Liao et al., 2016). In enzymatic reactions, a substrate is a molecule being converted by the enzyme catalyst. By binding with and acting on specific substrates, enzymes allow for dramatically accelerated rates in the transformation of their substrates (Bar-Even et al., 2011). Designing enzymes that can bind to specific substrates is a critical yet challenging problem.
@@
Recently, deep learning methods are promising in protein design (Huang et al., 2016; Pearce & Zhang, 2021). Existing deep learning approaches for functional protein design fall into three categories: (a) generative models for protein sequence design guided by fitness landscape (Brookes et al., 2019; Rives et al., 2021; Ren et al., 2022; Song & Li, 2023; Wang et al., 2023); (b) structure to sequence generation based on a targeted protein backbone (Ingraham et al., 2019; Dauparas et al., 2022; Zheng et al., 2023); (c) co-design of a backbone structure and a protein sequence that encodes it (Anishchenko et al., 2021; Wang et al., 2022; Shi et al., 2022; Yeh et al., 2023). However, these methods face many limitations on enzyme design. First, fitness-guided approaches are limited by the lack of fitness data for the vast majority of enzyme families. Second, the structures of many enzymes remain unknown (Binz et al., 2005; Fischman & Ofran, 2018). Third, prior approaches do not model substrates in enzyme design process. Finally, there is no unified model applicable to massive and diverse enzyme families.
@@
In this paper, we aim to develop a unified generative model to design functional enzymes across thousands of enzyme families. The key design idea lies in the notion that the biological function of an enzyme is enabled by a subset of residues (amino acids), known as functionally important sites (Bickel et al., 2002; Chakrabarti & Lanczycki, 2007; Wang et al., 2022), and that an ideal enzyme should be able to bind its substrate(s) in enzymatic reactions. Therefore, we formulate the enzyme design problem as jointly generating the enzyme sequence and backbone structure given an enzyme’s functional category, automatically mined functionally important sites, and associated substrates in the catalytic reactions of the enzyme.
@@
To this end, we propose EnzyGen, a unified model to co-design enzyme sequence and backbone structure. EnzyGen comprises an enzyme modeling module and a substrate representation module. We devise neighborhood attentive equivariant layers (NAELs) to jointly model enzyme amino acid sequence and backbone structure coordinates. Each NAEL consists of a global attention sub-layer and a neighborhood equivariant sub-layer. The global attention sub-layer captures correlations among all residues within an entire enzyme sequence, while the neighborhood equivariant sub-layer updates residue representations and coordinates based on nearest neighbors in the 3D space. This architectural design facilitates information exchange with varying levels of granularity, promoting a comprehensive interaction among different residues. The substrate representation module also uses neighborhood equivariant layers to model a given substrate. Furthermore, EnzyGen learns embeddings for enzyme tags corresponding to the BRENDA databaseto enable transfer among enzymes within the same family branch. To train EnzyGen, we design a joint training objective consisting of an amino acid type prediction loss, a residue coordinate reconstruction loss and an enzyme-substrate interaction loss.
@@
  • We propose EnzyGen to jointly generate enzyme sequence and backbone structure. EnzyGen is the first unified enzyme design model across thousands of enzyme families.
  • We create EnzyBench, a comprehensive benchmark for training and evaluating enzyme design. It includes $101,974$ PDB entries across $3,157$ enzyme families, and a set of three metrics to evaluate enzyme quality prior to wet-lab experiments. EnzyBench includes a test set with $323$ families for various function validation.
  • We conduct experiments to evaluate EnzyGen and existing functional protein design methods. EnzyGen outperforms the best prior baseline by 10.79% in terms of substrate binding affinity. We show that enzymes designed by EnzyGen exhibit the best average enzyme-substrate ESP score (Kroll et al., 2023) of 0.65, substrate binding affinity of -9.44, and AlphaFold2 pLDDT (Jumper et al., 2021) of 87.45 across 323 families (BRENDA enzyme classification fourth-level categories).
@@ 2 Related Work
Methods for Functional Protein Design. Functional protein design has been studied with a wide variety of methods. Some works focus on designing functional proteins guided by fitness landscape, including searching (Brookes & Listgarten, 2018; Brookes et al., 2019; Kumar & Levine, 2020; Das et al., 2021; Hoffman et al., 2022; Melnyk et al., 2021; Anishchenko et al., 2021; Ren et al., 2022) or directly generating (Jain et al., 2022; Song & Li, 2023) protein sequences through deep generative models. Another category of methods concentrates on designing protein sequences capable of folding into backbone structures to fulfill specific functions (Wang et al., 2018; Strokach et al., 2020; Dauparas et al., 2022; Hsu et al., 2022; Sumida et al., 2023). Owing to the availability of large-scale data, some models are trained on extensive protein sequences from diverse families (Ferruz et al., 2022) or incorporate additional control tags for enhanced modeling (Madani et al., 2023). On the path from sequence to function, the role of structure emerges as a crucial intermediary. Consequently, some studies focus on the design of functional structures (Trippe et al., 2022; Watson et al., 2023). Recognizing the inherently complex interplay between protein sequence and structure, another set of works relies on the simultaneous generation of a new backbone structure and an amino acid sequence that folds into it (Anishchenko et al., 2021; Wang et al., 2022; Shi et al., 2022; Yeh et al., 2023; Song et al., 2023). Compared to the design of other functional proteins, functional enzyme design is largely underdeveloped. Some of the above methods have been applied to specific enzyme families, such as generating satisfactory enzyme sequences using deep generative models (Detlefsen et al., 2022; Lin et al., 2022a; Giessel et al., 2022). However, no previous model is capable of generating enzymes in all families (over 3,000).
@@ 3 Proposed Method: EnzyGen
An enzyme consists of a chain of amino acids (also called residues) connected by peptide bonds, which folds into a proper 3D structure. Let $\mathcal{A}$ be the set of 20 common amino acids. We denote the sequence of a $N$-residue enzyme by $\boldsymbol{s}=\{s_{1},s_{2},...,s_{N}\}\in\mathcal{A}^{N}$ and their $C_{\alpha}$ coordinates by $\boldsymbol{x}=[\boldsymbol{x}_{1},\boldsymbol{x}_{2},...,\boldsymbol{x}_{N}]^{T}\in\mathbb{R}^{N\times 3}$. For residue type $s_{i}\in\mathcal{A}$, where $i\in\{1,2,...,N\}$, we denote its one-hot encoding as $\boldsymbol{s}_{i}=\text{onehot}(s_{i})$ and the functionally important site index set as $\mathcal{M}$. We denote a substrate as $\mathcal{V}=\{v_{1},...,v_{m}\}$, where $m$ is the number of its atoms, and $v_{j}=(\boldsymbol{h}_{j},\boldsymbol{x}_{j})$ denotes an atom where $j\in\{1,2,...,m\}$. Here, $\boldsymbol{h}_{j}$ represents pre-computed chemical features, and $\boldsymbol{x}_{j}\in\mathbb{R}^{3}$ is the corresponding coordinate. Each substrate is paired with a binary label $y$ indicating whether the enzyme can bind with it. Inspired by multilingual machine translation models in natural language processing, which employ language tags to guide model generation in specific target languages (Johnson et al., 2017; Liu et al., 2020), we similarly use enzyme family tags to facilitate enzyme design with desired functions. To provide a systematical and professional enzyme family classification, we leverage the enzyme classification (EC) tree in the BRENDA database (Schomburg et al., 2004). For example, the EC tag 1.1.1.1 denotes enzyme family of alcohol dehydrogenase.
The problem studied in this paper can be formulated as follows: given four-level EC tag $c=\{c_{1},c_{2},c_{3},c_{4}\}$, a substrate $\mathcal{V}$, a functionally important site index set $\mathcal{M}$, its residue set $\boldsymbol{s}_{\mathcal{M}}$ and the corresponding 3D coordinates $\boldsymbol{x}_{\mathcal{M}}$, generate an enzyme sequence $\boldsymbol{s}$ and $C_{\alpha}$ coordinates $\boldsymbol{x}$ of $N$ residues satisfying the substrate binding constraint $y$. Essentially, we aim to learn a generative model with probability $P(\boldsymbol{s},\boldsymbol{x},y|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c,\mathcal{V})$. At inference time, we only use EC tag $c$, the functionally important sites $\boldsymbol{s}_{M}$ and $\boldsymbol{x}_{M}$ to generate an enzyme.
@@ 3.1 Overall Model Architecture
$\footnotesize\begin{split}P(\boldsymbol{s},\boldsymbol{x},y&|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c,\mathcal{V};\theta)=P(\boldsymbol{s},\boldsymbol{x}|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)\\ &\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\cdot P(y|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c,\mathcal{V};\theta)\\ P(\boldsymbol{s},\boldsymbol{x}&|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)=P(\boldsymbol{s}|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)\\ &\hskip 17.00024pt\hskip 17.00024pt\hskip 17.00024pt\qquad\cdot P(\boldsymbol{x}|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)\\ P(\boldsymbol{s}&|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)=\Pi_{i=1\&i\notin\mathcal{M}}^{N}P(\boldsymbol{s}_{i}|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)\\ P(\boldsymbol{s}_{i}&|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)=\mathrm{Softmax}(W_{\mathcal{A}}\cdot\boldsymbol{h}_{i}^{L_{e}})\\ P(\boldsymbol{x}&|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)=\Pi_{i=1\&i\notin\mathcal{M}}^{N}P(\boldsymbol{x}_{i}|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}},c;\theta)\\ \boldsymbol{x}_{i}&|\boldsymbol{s}_{\mathcal{M}},\boldsymbol{x}_{\mathcal{M}}\sim\mathcal{N}(\boldsymbol{x}_{i}^{L_{e}};\lambda I)\end{split}$ (1)
where $W_{\mathcal{A}}$ is the embedding matrix for 20 common amino acids, $\boldsymbol{h}_{i}^{L_{e}}$ is the output embedding for $i^{th}$ residue at the last layer of enzyme modeling module and $\boldsymbol{x}_{i}^{L_{e}}$ is the corresponding output coordinate. $\mathcal{N}(\boldsymbol{x}_{i}^{L_{e}};\lambda I)$ is the Gaussian distribution with mean $\boldsymbol{x}_{i}^{L_{e}}$ and covariance matrix $\lambda I$ (II is the identity matrix). $\lambda$ is a hyperparameter. To find the optimal $\theta$, we maximize the conditional log likelihood (i.e., minimizing the negative log likelihood):
@@
[…]
@@
[…]